OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point:(2,-6) ; Slope:-3/4

OpenStudy (anonymous):

what r ur options?

OpenStudy (anonymous):

A.y-6=-3/4(x+2) b.y-6=3/4(x-2) c.y+6=-3/4(x-2) d.y+6=-3/4(x+2)

OpenStudy (anonymous):

which do u think it is?

OpenStudy (anonymous):

i just realised that i put this in history and not mathematics....

OpenStudy (anonymous):

its fine i can still help u

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

you r CORRECT!!!!!!!!!!!!! :)

OpenStudy (anonymous):

great job

OpenStudy (anonymous):

yaaay thank you do you mind helping me with a couple more?

OpenStudy (anonymous):

no i dont mind